OpenStudy (anonymous):

Medal for best answers! 1. A dozen apples cost $2.55. At this rate, how much would 8 apples cost? A. $20.40 B. $3.83 C. $0.21 D. $1.70 2. Solve the proportion. Where necessary, round to the nearest hundredth. 8/2 = x + 3/4 A. 13 B. 1 C. 7 D. 16

OpenStudy (bibby):

for 1, find the price of each individual apple for 2 cross multiply

OpenStudy (anonymous):

Ok, thanks

OpenStudy (anonymous):

@bibby 1. is D, am I correct?

OpenStudy (bibby):

let's see. 12apples=2.55 divide both sides by 12 \(\huge \frac{1}{\cancel{12}}\cancel{12}x=2.55*\frac{1}{12}\) 0.2125 for each apple. 8 apples costs 0.2125*8 or 1.70

OpenStudy (bibby):

yeah, good job

OpenStudy (anonymous):

1. $1.70

OpenStudy (bibby):

I guess you could technically do that in a proportion too. \(\huge \frac{2.55}{12}=\frac{x}{8}\) 20.4=12x x=1.7

OpenStudy (anonymous):

i no